From Ciswikidb
Jump to: navigation, search

Meeting Information


To connect to the meeting, please do the following: Time: Wednesday at 9:00 AM EST.

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/984025598.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 984025598.
  3. To join via phone, use one of the following numbers and the Conference ID: 984025598.
    • Dial Either +1 408 740 7256 or +1 888 240 2560 for US or Canada



Long Term Schedule: Gantt (Outstanding tasks)

  1. 7/26/21 Start Hot Check-out

New items for discussion:

Short term schedule update:
This week:

Reports by Group: CIS/EGG:

  1. Gun/Depo chamber bake started and completed
  2. Gun HVPS tank topped to 60 psig SF6; NF3 charged as well
  3. R30 resistor installed for gun HV processing
  4. Gun anode picoammeter connected
  5. Decarad probes placed for gun
  6. MDSK201 interlock in gun HVPS bypassed for HV commissioning (See ABIL5170)
  7. Steel puck in place
  8. Gun HV Processing under way (UITFList 1663); up to 182kV so far
  9. Gun rebuild schedule

DC Power:

  1. Corrector installation & cable reroute mostly complete, turned over to DC power
  2. Correctors inventoried following some naming confusion (MHB vs MBH); MBH = Magnet Big Haimson (2.75" bore); MHB family will be renamed MLH = Magnet Little Haimson (1.5" bore)
  3. Chris Gould making nomenclature changes / Gary Croke & Michele Joyce making controls changes / Eric Forman to update Quick Reference Drawing / Shaun to take care of song sheets
  4. Changes made in UED workspace for names and field maps; ready to merge soon. (Thanks Gary!)
  5. Raster installation - completed; per Bill Gunning. (Thanks!) Also, thank you for software and channel access integration by Weiwei & Brian and John for staying on top of things!
  6. Solenoid PS being installed

Instrumentation and Controls:

  1. BPMs connections good - save iocipmm904 removal - iocipmm901 installed in its place
  2. BPM IOCs updated with new configuration. (Thanks Kyle!)
  3. Install target mechanism (Completed), motor controls tested - Survey required
  4. Insertables APP undated with new installation (Thanks Scott!); iocitf2 rebooted.
  5. BCM project software integration under way between Musson/Seaton and Sue Witherspoon
  6. Chris Norris to hook up viewers next week.
  7. Brian Freeman available for adViewer work for video camera for ITVM604A; some focusing still required.


  1. 6GHz > 1.5GHz Down-converter installed for bunch length system


  1. Spectrometer pumps on EPICs, Turbo disconnected


  1. BLM locations being set 7/21 following this meeting
  2. FSD inputs for magnet and BCM being prepared

ME / Installation:

  1. Nomenclature for elements disseminated and labelled
  2. Build 80/20 support structure for Pb (Completed)
  3. Ready to install Al base plates for FC3 region rework as well as solenoid stand. 80/20 reworked for support
  4. Pb stacked; RCG required modification due to permanent shielding req delay.
  5. New designs for FC3 and FC4 with a great many thanks to Shaun for the design work and John/Xi for Pb arrangement and table construction!
  6. Songsheet being updated by Shaun; be sure to let him know when changes are made!
  7. Status of permanent shielding purchase?


  1. ESAD, COO, and draft response to ERR under way by Gigi. ESAD/COO signed and turned over to Accelerator DSO.


Still to do:

  • Add Mu metal over bellows on wastewater line.

Next Week:

  • Mount bottle rack in Labyrinth (Staged); remove gas bottles to rack.
  • Unistrut mounting for cables - Neal will do when available
  • Engineering work on beam line starting

2 weeks:

  • Pull Heliax from Spectrum analyzer/Control room >> Buncher
  • 4th terminal system in Control room under assembly/rebuild
  • Items from To-do list...

Engineering Other business

Reports on Open Action Items:

Action Items:

Back to Main Meeting Minutes Page